OpenStudy (anonymous):

Simplify the following expression: (5x^2 + 3x + 4) − (2x^2 + 5x − 1). If the final answer is written in the form Ax2 + Bx + C, what is the value of A?

OpenStudy (anonymous):

(5x2 + 3x + 4) − (2x2 + 5x − 1) = 5x2 + 3x + 4 − 2x2 - 5x + 1 = 5x2 - 2x2 + 3x - 5x + 4 + 1 = 3x2 - 2x + 5. A = 3, B = -2, C = 5 is what i got, but is A = 3, B = -2, C = 5 the standard form ?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

okay well where i have to put my answer, i have to write it in. and its very specific. even if i get the answer right, if i dont write i exactly how the computer has it, ill get it wrong. so what should i type in ????

OpenStudy (anonymous):

is it on flvs?

OpenStudy (anonymous):

yes. what should i put in ?

OpenStudy (anonymous):

i dont see how that matters though.

OpenStudy (anonymous):

the answer then is 3....i say that cause im on flvs also and gotten used to how it wants the answers

OpenStudy (anonymous):

why would i jus write 3 ??

OpenStudy (anonymous):

answer for blank:[__________] is what the thing looks like right?.....if so then it just wants you answer number only

OpenStudy (anonymous):

okay then. i think its wants more then jus 3 but fine

OpenStudy (anonymous):

Simplify the following expression: (x + 6y) - (3x − 10y). If the final answer is written in the form Ax + By, what is the value of A?

OpenStudy (anonymous):

same thing

OpenStudy (anonymous):

sorry 1 sec

OpenStudy (anonymous):

sorry my neighbor was being stuborn and nocking

OpenStudy (anonymous):

its okay. so what would i put in for this one

OpenStudy (anonymous):

16y-2x A=16 B= -2 so i would put 16

OpenStudy (anonymous):

i got that one wrong.

OpenStudy (anonymous):

was it one i answer?

OpenStudy (anonymous):

yeaa the last one

OpenStudy (anonymous):

hmm well im sorry i got that 1 wrong

OpenStudy (anonymous):

its okay. can u help with a few more ?

OpenStudy (anonymous):

if i understand them sure

OpenStudy (anonymous):

Choose the correct simplification of the expression (d^3)^5. d^8 d^15 d^243 d

OpenStudy (anonymous):

ima think its d^15 but let me check my notes i did that earlier

OpenStudy (anonymous):

yeaa it is ^_^ Choose the correct simplification of the expression (x^2y)^2. x^4y^3 x^6y^6 x^4y^2 xy^4

OpenStudy (anonymous):

c x^4y^2

OpenStudy (anonymous):

Compare and Contrast: Below are two expressions. Simplify each and choose the statement that is true about each. Expression #1 (4x)2(x) Expression #2 (6x2)4 The exponents in Expression #1 are greater than the exponents of Expression #2. The exponents on Expression #2 are greater than the exponents of Expression #1. The exponents of Expression #1 are the same as the exponents of Expression #2. The relationship cannot be determined with the given information.

OpenStudy (anonymous):

essay question?

OpenStudy (anonymous):

no. multiple choice

OpenStudy (anonymous):

ah ok 1 sec

OpenStudy (anonymous):

Compare and Contrast: Below are two expressions. Simplify each and choose the statement that is true about each. Expression #1 (4x)2(x) Expression #2 (6x2)4

OpenStudy (anonymous):

The exponents in Expression #1 are greater than the exponents of Expression #2. The exponents on Expression #2 are greater than the exponents of Expression #1. The exponents of Expression #1 are the same as the exponents of Expression #2. The relationship cannot be determined with the given information.

OpenStudy (anonymous):

i would say answer b

OpenStudy (anonymous):

oh okay Choose the correct simplification of the expression a^2 • a^3. a a^6 a^5 a^−1

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Compare and Contrast: Below are two expressions. Simplify each and then choose the statement that is true. Expression #1 (3x2)3x2 Expression #2 (3x3)2(x2) The exponents in Expression #1 are greater than the exponents of Expression #2. The exponents on Expression #2 are greater than the exponents of Expression #1. The exponents of Expression #1 are the same as the exponents of Expression #2. The relationship cannot be determined with the given information.

OpenStudy (anonymous):

i would say b agin....the compare and contrast 1s i say for you to double check my answers

OpenStudy (anonymous):

okay ill look over it. Choose the correct simplification of the expression c^2 • c^9. c^11 c^7 c^18 c

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Choose the correct simplification of the expression (a^2)^3. a^5 a^8 a a^6

OpenStudy (anonymous):

d

OpenStudy (anonymous):

Choose the correct simplification of the expression (3x)^4. 81x^4 12x^5 81x^5 12x^4

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x^3y^4)(2x^2y^6)

OpenStudy (anonymous):

3*2=6 x^3*x^2=x^5 y^4*y^6=y^10 6x^5y^10 you can double check if you'd like to make sure im correct but the complete senteces is all you sorry

OpenStudy (anonymous):

no problem. thanks !

OpenStudy (anonymous):

glad to help

OpenStudy (anonymous):

got them all right ^_^